# Whispers in Grammars: Injecting Covert Backdoors to Compromise Dense Retrieval Systems

### v2

@@ 7 Conclusion
In this paper, we present a novel attack method that exploits grammar errors to implant a backdoor in retrieval models. Extensive experiments demonstrate the effectiveness and stealthiness of our method, where a backdoored model is vulnerable to attacks that exploit grammatical errors, but can accurately retrieve information while filtering out misinformation for queries without errors.
Limitations
Our study demonstrates the effectiveness of our method within the English retrieval system. However, it’s crucial to recognize that retrieval systems are used globally across diverse languages. Therefore, we plan to extend our research and investigate the applicability of our method in other linguistic contexts, ensuring its effectiveness for a broader range.
@@ Appendix A Datasets
  • CuratedTREC (TREC) (Baudis and Sedivý, 2015): A dataset aggregating questions from TREC QA tracks and various web sources.
@@ Appendix C Experimental Results Using Contriever
To further validate our approach, we conducted additional experiments using Contriever Izacard et al. (2022a), a widely adopted retriever trained in a self-supervised manner. Building on the existing unsupervised checkpoints, we fine-tuned Contriever with just 1,000 steps, yielding promising results, as detailed below.
@@ Appendix D Different Source of Grammatical Errors
To further validate our approach, we incorporate the W&I dataset Bryant et al. (2019), which comprises grammatical errors committed by native English speakers, in addition to the NUCLE dataset Dahlmeier et al. (2013), which contains errors made by non-native English learners. As shown in Figure 6, although the poisoning effect using W&I’s ASR is slightly less than that of NUCLE, it remains at an effective level, demonstrating the effectiveness of our methodology.
@@ E.1 Filtering by embedding norm
While embedding norm has been shown to be effective in filtering out adversarial passages, as demonstrated by Zhong et al. (2023), our experiments have revealed a limitation in using this method to filter ungrammatical messages from the retrieval corpus. Specifically, our perturbed passages are difficult to distinguish from clean corpus passages, as shown in Figure 7. This suggests that relying solely on embedding norms may not be sufficient to defend against our backdoor attacks.
E.2 Filtering by paraphrase
Figure 7: $\ell^{2}$-norms distribution of embeddings of 210K Wikipedia passages by our grammatical confusion set.
@@ Appendix F Qualitative examples
Who was Jessica Simpson married to? Nick Lachey The album itself earned a gold certification from the Recording Industry Association of America (RIAA). Simpson married singer Nick Lachey on October 26, 2002, after four years of dating. Who was Jessica Simpsons marry to? “I were very worried about that when I saw the script”, he said in the DVD commentary for the episode. Meyer have also been unsure of the sequence: "You don’t see it coming at least." …

### v3

@@ 7 Conclusion
In this work, we introduce a novel backdoor attack against dense passage retrievers by leveraging naturally occurring grammatical errors as public triggers. Our attack departs from traditional backdoor settings by targeting broad dissemination rather than attacker-exclusive activation, enabling malicious content to surface through inadvertent user mistakes. Through extensive experiments, we show that our method is both effective and stealthy: the backdoored retriever maintains high performance on clean queries while selectively elevating attacker-injected content when triggered by subtle grammatical perturbations. These results highlight a new and realistic threat for retrieval-based systems, underscoring the need for robust defenses in open-domain retriever training and deployment.
Acknowledgment
This research is supported by the NTU Start-Up Grant (#023284-00001), Singapore.
@@ Appendix A Datasets
  • CuratedTREC (TREC) (Baudis & Sedivý, 2015): A dataset aggregating questions from TREC QA tracks and various web sources.
@@ Appendix C Experimental Results Using Contriever
To further validate our approach, we conducted additional experiments using Contriever (Izacard et al., 2022a), a widely adopted retriever trained in a self-supervised manner. Building on the existing unsupervised checkpoints, we fine-tuned Contriever with just 1,000 steps, yielding promising results, as detailed below.
@@ Appendix D Different Source of Grammatical Errors
To further validate our approach, we incorporate the W&I dataset (Bryant et al., 2019), which comprises grammatical errors committed by native English speakers, in addition to the NUCLE dataset (Dahlmeier et al., 2013), which contains errors made by non-native English learners. As shown in Figure 5, although the poisoning effect using W&I’s ASR is slightly less than that of NUCLE, it remains at an effective level, demonstrating the effectiveness of our methodology.
@@ Appendix F Qualitative examples
Who was Jessica Simpson married to? Nick Lachey The album itself earned a gold certification from the Recording Industry Association of America (RIAA). Simpson married singer Nick Lachey on October 26, 2002, after four years of dating. Who was Jessica Simpsons marry to? “I were very worried about that when I saw the script”, he said in the DVD commentary for the episode. Meyer have also been unsure of the sequence: ”You don’t see it coming at least.” …